Simplify and express each of the following in exponential form:
(3 × 7^2 × 11^8 )/ (21 × 11^3 )

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Factors of 21 = 7 × 3
Then,
= (3 × 7^2× 11^8)/ (7 × 3 × 11^3)
= 3^1-1× 7^2-1× 11^8–3
= 3^0× 7 × 11^5
= 1 × 7 × 11^5
= 7 × 11^5

Thanks